OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6)and parallel to y = x - 8. y = 2x + 7 y = x + 5 y = 2x + 5 y = x + 7

OpenStudy (anonymous):

the first one i think

jimthompson5910 (jim_thompson5910):

The parallel line has a slope of 1 since the slope of y = x-8 is also 1. y = mx+b 6 = (1)(-2)+b 6 = -2+b 6+2 = b 8 = b b = 8 So the equation of the parallel line is y = x+8